The probability that an American CEO can transact business in a foreign language is .40. Fourteen American CEOs are chosen at random.

  

(a)

What is the probability that none can transact business in a foreign language? (Round your answer to 4 decimal places.)

  

  Probability   

  

(b)

What is the probability that at least two can transact business in a foreign language? (Round your answer to 4 decimal places.)

  

  Probability   

  

(c)

What is the probability that all 14 can transact business in a foreign language? (Round your answer to 10 decimal places.)

  

  Probability   
Answer #1

a)probability that none can transact business in a foreign  language =(0.60)14 =0.0008

b)

probability at least two can transact business in a foreign language:

probability = P(X>=2)= 1-P(X<=1)= 1-∑x=0x-1     (nCx)px(1−p)(n-x)    = 0.9919

c)

  probability that all 14 can transact business in a foreign language =0.414 =0.0000026844